OpenStudy (anonymous):

(4^2x)+(4^x)-6=0 Solve the equation. Express the solution set in terms of natural logarithms. then use a calculator to obtain a decimal approximation to two decimal places.

OpenStudy (tkhunny):

How would you solve this one? \(w^{2} + w - 6 = 0\)

OpenStudy (anonymous):

factor?

OpenStudy (tkhunny):

Let's see.

OpenStudy (anonymous):

-3 and 2

OpenStudy (anonymous):

but it has to be in terms of Ln

OpenStudy (tkhunny):

It was more the process I was hoping to see. Your work is far more valuable than the result. \(w^{2}+w−6=0\) \((w+3)(w-2) = 0\) \(w = -3\;or\;w = 2\) Perfect. Now, do EXACTLY the same thing with your original equation. \((4^{2x})+(4^x)-6=0\) \((4^{x})^{2}+(4^x)-6=0\) \((4^{x} + 3)(4^{x} - 2) = 0\) Are you seeing it?

OpenStudy (anonymous):

yes..

OpenStudy (tkhunny):

Always look for Quadratic Forms. \(Something^{2} + a\cdot Something + b = 0\) This is why you did WAY MORE factoring problems than you thought was necessary back in Algebra. Good work.

OpenStudy (anonymous):

but i have to write it in terms of Ln, thats what i dont understand

OpenStudy (anonymous):

\[ w=4^x\implies \ln w= x\ln4\implies x=\frac{\ln w}{\ln 4} \]

hartnn (hartnn):

did that help you further ? which step are u still stuck on ?

OpenStudy (anonymous):

The correct answer is ln2/ln4 how is that?

hartnn (hartnn):

didn't you get w =2 ?? and w=-3 is dicarded, because log og negative number is not defined.

hartnn (hartnn):

log of**

hartnn (hartnn):

\(w=4^x\implies \ln w= x\ln4\implies x=\frac{\ln w}{\ln 4} =\frac{\ln 2}{\ln 4} \)

OpenStudy (anonymous):

ok that makes sense now